# Mind the Confidence Gap: Overconfidence, Calibration, and Distractor Effects in Large Language Models

### v2

@@ 4.3 Influence of Fine-Tuning Regime and Model Architecture
(a) ECE comparison across question types
(b) Date
(c) Number
(d) Person
(e) Place
4.4 Effect of Model Size within LLM Families
We disentangle parameter count from other factors by comparing the smallest, mid-sized, and largest checkpoints released by each provider. In the normal setting ($\mathcal{N}$), accuracy increases monotonically with scale for both families; however, calibration improves much faster for the OpenAI series. GPT-4o already attains an ECE of 0.450 on SimpleQA. This suggests that RLHF alignment, used uniformly across GPT-4 models, amplifies the natural size-driven gains in self-assessment that emerge from scaling alone.
Introducing distractor options ($\mathcal{D}$) radically alters the picture. The largest relative accuracy jumps occur in the smallest models as shown in Figure 3: GPT-4o-mini rockets from 8.5% to 47.4% accuracy on SimpleQA (+461%), and LLaMA-3-8b leaps +819% over the same split. In contrast, their flagship counterparts—GPT-4o and LLaMA-70b—gain a more modest +109% and +338%, respectively. Yet these headline boosts do not translate into equally dramatic calibration improvements. After distractors, GPT-4o compresses its ECE by 92% (to 0.037), whereas GPT-4o-mini still lingers above 0.32. A mirror pattern holds for LLaMA: the 70B model cuts ECE by 69%, finishing at 0.24, while the 8B base remains mis-calibrated (ECE 0.36) despite its vast accuracy lift. When we extend this analysis to FaVIQ and TriviaQA, the same ranking by model size holds but with attenuated returns. On FaVIQ, distractors yield solid—but more moderate—accuracy boosts across all scales. On TriviaQA, where the baseline performance is already high, relative gains shrink into the low-teens, indicating only marginal benefit from explicit distractors.
Two key takeaways emerge from our analysis. First, small models acquire factual knowledge more quickly than they learn to assess their own confidence: adding explicit answer choices boosts accuracy but yields poorly calibrated probability estimates. Second, increasing model scale primarily improves a model’s ability to quantify its certainty rather than uncover new knowledge. Consequently, when downstream tasks rely directly on confidence scores—such as risk-aware planning or answer adjudication—larger models offer more trustworthy probabilities. In contrast, in settings where latency or compute cost is the main constraint and some post-hoc temperature scaling is acceptable, smaller models can still deliver adequate confidence estimates so long as they’re provided with distractors or retrieval-augmented context.
4.5 Performance Across Question Types
The SimpleQA dataset already categorizes 4326 questions into four non-overlapping types—Date (1418), Number (663), Person (1041), and Place (427)—and we evaluate each model’s accuracy and ECE in both the free-generation ($\mathcal{N}$) and distractor-augmented ($\mathcal{D}$) settings. To better understand calibration weaknesses, we analyze model performance and confidence alignment across these question types (Figure 4). Person-based queries are most challenging, likely due to name ambiguities and inherent variability in names, overlapping roles, and contextual dependencies that require deeper reasoning beyond surface-level pattern matching. LLMs frequently confuse historical figures with similar names, but providing structured answer choices significantly improves accuracy, suggesting that explicit disambiguation helps mitigate uncertainty in this category. In contrast, place-based queries exhibit relatively strong performance across both settings, indicating that geographic knowledge is well-represented in pretraining. However, calibration improvements vary (Figure 4(a)): the person category sees highest relative ECE drop (69%), while place category shows the lowest (49%). This suggests that structured choices help correct overconfidence in ambiguous queries but offer limited calibration gains when models already retrieve knowledge with high confidence. These results demonstrate that miscalibration depends on both task framing and knowledge representation, not just model scale or architecture. While structured reasoning improves confidence alignment in person-based queries, factual retrieval tasks like place-based questions may require alternative calibration strategies to prevent persistent overconfidence.
@@ 5 Conclusion
Limitations
While our work breaks new ground in structured distractor–augmented calibration, we note three main limitations:
Model & task scope.
We cover nine state-of-the-art LLMs across three single-turn factoid QA benchmarks, which gives a broad but still focused view of calibration behavior. Extending to domain-specialized models, multi-hop reasoning, and generative tasks is an important direction for future validation.
Distractor realism.
To ensure consistency, all distractors were produced by a controlled LLM prompt. Although this may not capture every nuance of real-world retrieval, we followed strict prompting guidelines and shuffled options to minimize artefacts Zheng et al. (2023)—nevertheless, evaluating with human-curated or retrieval-sourced distractors remains valuable future work.
Automated-judge bias.
We use GPT-4o-mini as our primary correctness judge, complemented by a human evaluation (Cohen’s $\kappa$ = 0.82, see Appendix B for details) to confirm its reliability. Yet, subtle misclassifications could still subtly affect accuracy metrics, so applying complementary human or ensemble-based judges would further strengthen the results.
@@ Appendix A Custom Prompts
60 "answer": "Your answer here",
61 "confidence_score": number between 0-100
@@
89 "answer": "Your answer here",
90 "confidence_score": number between 0-100
@@ 85. The answers should be realistic and not obviously wrong
65- The output should be a single JSON object with key "wrong_answers"
@@ Appendix B Same LLM judge as Prediction LLM
We employ the same LLM as both the judge and the base model responsible for predicting answers to the questions. The performance metrics are presented in Table 2. Upon manually inspecting instances from smaller LLMs, we observe that the LLM judge occasionally misclassifies responses and refrains from assigning NOT_ATTEMPTED to certain data points. This can be seen by comparing $\mathcal{N}_{none}$ in Table1 and 2 for the SimpleQA dataset. To address this issue and ensure consistency, we use GPT-4o-mini as the LLM judge across all models. We also create reliability diagrams of pipeline where LLM-judge was different and the graphs are shown in Figure 5.
@@
Appendix C Distractor Generation
Our distractors were generated using GPT-4o-mini with a carefully designed prompt ensuring distractors were factually incorrect yet plausible (matching answer type, specificity, and context). To validate distractor plausibility further, we conducted a manual inspection on a subset (500+ randomly examples) from SimpleQA dataset, confirming that distractors were consistently plausible & contextually relevant.
Appendix D LLMs details
Table D surveys nine leading LLMs, revealing how differences in scale, architecture, and pretraining data shape their alignment pipelines. The GPT-4 family and LLaMA-3 series are dense models trained on massive multimodal or text-only corpora, each undergoing both SFT and RLHF before instruct-tuning. In contrast, LLaMA-4-Scout employs a 16-expert MoE design over mixed vision-text data, while Gemma2-9b-it and Qwen-qwq-32b explore distillation-based and pure SFT regimes, respectively. Despite these varied strategies, all nine models receive dedicated fine-tuning and instruction-tuning to optimize performance and calibration in downstream QA and conversational tasks.
GPT-4o
(undisclosed / 128K)
GPT-4o-mini
(undisclosed / 128K)
Dense Multi-modal (web text, code, images) SFT + RLHF
GPT-4-turbo
(undisclosed / 128K)
Dense Multi-modal (web text, code, images) Pre-training + SFT + RLHF
LLaMA-3.1-8B-Instant
(8B / 128K)
LLaMA-3-8B-Instruct
(8B / 8K)
LLaMA-3-70B-Instruct
(70B / 8K)
LLaMA-4-Scout-17B
(17B / 10M)
Gemma2-9B-it
(9B / 8K)
Qwen-qwq-32B
(32B / 131K)
Appendix E Reliability Diagrams of Selected Models on FaVIQ and TriviaQA dataset

### v3

@@ 4.3 Influence of Fine-Tuning Regime and Model Architecture
4.4 Effect of Model Size within LLM Families
We disentangle parameter count from other factors by comparing the smallest, mid‑sized, and largest checkpoints released by each provider. In the normal setting ($\mathcal{N}$), accuracy increases monotonically with scale for both families; however, calibration improves much faster for the OpenAI series. GPT‑4o already attains an ECE of 0.450 on SimpleQA. This suggests that RLHF alignment, used uniformly across GPT‑4 models, amplifies the natural size‑driven gains in self‑assessment that emerge from scaling alone.
Introducing distractor options ($\mathcal{D}$) radically alters the picture. The largest relative accuracy jumps occur in the smallest models as shown in Figure 3: GPT‑4o‑mini rockets from 8.5% to 47.4% accuracy on SimpleQA (+461%), and LLaMA‑3‑8b leaps +819% over the same split. In contrast, their flagship counterparts—GPT‑4o and LLaMA‑70b—gain a more modest +109% and +338%, respectively. Yet these headline boosts do not translate into equally dramatic calibration improvements. After distractors, GPT‑4o compresses its ECE by 92% (to 0.037), whereas GPT‑4o‑mini still lingers above 0.32. A mirror pattern holds for LLaMA: the 70B model cuts ECE by 69%, finishing at 0.24, while the 8B base remains mis‑calibrated (ECE 0.36) despite its vast accuracy lift. When we extend this analysis to FaVIQ and TriviaQA, the same ranking by model size holds but with attenuated returns. On FaVIQ, distractors yield solid—but more moderate—accuracy boosts across all scales. On TriviaQA, where the baseline performance is already high, relative gains shrink into the low-teens, indicating only marginal benefit from explicit distractors.
(a) Date
(b) Number
(c) Person
(d) Place
Two key takeaways emerge from our analysis. First, small models acquire factual knowledge more quickly than they learn to assess their own confidence: adding explicit answer choices boosts accuracy but yields poorly calibrated probability estimates. Second, increasing model scale primarily improves a model’s ability to quantify its certainty rather than uncover new knowledge. Consequently, when downstream tasks rely directly on confidence scores—such as risk‐aware planning or answer adjudication—larger models offer more trustworthy probabilities. In contrast, in settings where latency or compute cost is the main constraint and some post‐hoc temperature scaling is acceptable, smaller models can still deliver adequate confidence estimates so long as they’re provided with distractors or retrieval‐augmented context.
4.5 Performance Across Question Types
The SimpleQA dataset already categorizes 4326 questions into four non-overlapping types—Date (1418), Number (663), Person (1041), and Place (427)—and we evaluate each model’s accuracy and ECE in both the free-generation ($\mathcal{N}$) and distractor-augmented ($\mathcal{D}$) settings. To better understand calibration weaknesses, we analyze model performance and confidence alignment across these question types (Figure 4). Person-based queries are most challenging, likely due to name ambiguities and inherent variability in names, overlapping roles, and contextual dependencies that require deeper reasoning beyond surface-level pattern matching. LLMs frequently confuse historical figures with similar names, but providing structured answer choices significantly improves accuracy, suggesting that explicit disambiguation helps mitigate uncertainty in this category. In contrast, place-based queries exhibit relatively strong performance across both settings, indicating that geographic knowledge is well-represented in pretraining. However, calibration improvements vary (Table 3): the person category sees highest relative ECE drop (69%), while place category shows the lowest (49%). This suggests that structured choices help correct overconfidence in ambiguous queries but offer limited calibration gains when models already retrieve knowledge with high confidence. These results demonstrate that miscalibration depends on both task framing and knowledge representation, not just model scale or architecture. While structured reasoning improves confidence alignment in person-based queries, factual retrieval tasks like place-based questions may require alternative calibration strategies to prevent persistent overconfidence.
Table 3: ECE comparison across question types.
@@ 5 Conclusion
6 Limitations
Generator/Judge Dependence.
Our distractor-augmented setting fixes a single generator and a single judge (GPT-4o-mini) for consistency across models. While this reduces rubric drift and style confounds, it also risks bias from an “AI-generates/AI-judges” loop. We mitigate subjectivity via type- and format-matched distractor prompts, overlap and plausibility checks, and human spot-checks by three reviewers with disagreement resolution. Nevertheless, results should be interpreted as conditional on this (generator, judge) pair. We release prompts to support replication with alternative generators and judges. Exploring cross-model distractor sources and independent judges (including human-only adjudication) is important future work; our present study prioritizes comparability under a fixed setup and does not claim generator/judge invariance.
@@ Appendix A Custom Prompts
60 ‘answer’: ‘Your answer here’,
61 ‘confidence_score’: number between 0-100
@@
89 ‘answer’: ‘Your answer here’,
90 ‘confidence_score’: number between 0-100
@@ 85. The answers should be realistic and not obviously wrong
65- The output should be a single JSON object with key ‘wrong_answers’
@@ Appendix B Same LLM judge as Prediction LLM
We employ the same LLM as both the judge and the base model responsible for predicting answers to the questions. The performance metrics are presented in Table 4. Upon manually inspecting instances from smaller LLMs, we observe that the LLM judge occasionally misclassifies responses and refrains from assigning NOT_ATTEMPTED to certain data points. This can be seen by comparing $\mathcal{N}_{none}$ in Table2 and 4 for the SimpleQA dataset. To address this issue and ensure consistency, we use GPT-4o-mini as the LLM judge across all models. We also create reliability diagrams of pipeline where LLM-judge was different and the graphs are shown in Figure 5.
@@
Appendix C Reliability Diagrams of Selected Models on FaVIQ and TriviaQA dataset